A random sample of n=49 observations is drawn from a population with a mean equal to15

and a standard deviation equal to 14.

a.

Give the mean and standard deviation of the​ (repeated) sampling distribution

x overbarx.

b.

Describe the shape of the sampling distribution of

x overbarx. Does your answer depend on the sample​ size?

c.

Calculate the standard normal​ z-score corresponding to a value of

x overbarxequals=12.5.

d.

Calculate the standard normal​ z-score corresponding to a value of

x overbarxequals=19.

a.

mu Subscript x overbarμxequals=1515

sigma Subscript x overbarσxequals=22

b. Choose the correct answer below.

A.

The shape is that of a normal distribution and depends on the sample size.

Your answer is correct.

B.

The shape is that of a uniform distribution and depends on the sample size.

C.

The shape is that of a uniform distribution and​ doesn't depend on the sample size.

D.

The shape is that of a normal distribution and​ doesn't depend on the sample size.

c.

z=___________

d.

z=___________

Answers a and b are correct! Need the correct answer for c and d.
Answer #1

c)

as z score =(X-mean)/std error

hence z-score =(12.5-15)/2=-1.25

d)

z-score =(19-15)/2=2